# A kernel conditional two-sample test

### v1

@@ D.4 Main result
Theorem D.23.
Under the notations of Section D.1, let $R\in\mathcal{L}_{\mathrm{b}}(\mathcal{G})$ be self-adjoint, positive semi-definite, and $S\in\mathbb{R}_{>0}$. Assume that $\lVert h^{\star}\rVert_{\mathcal{H}}\leq S$, and assume one of the following and define $\beta_{\lambda}$ and $\mathcal{N}$ accordingly:
@@
    Assumptions D.5 and D.13 hold, $R=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$ for some $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, and $\eta$ is $R$-subgaussian conditionally on $\mathcal{F}$. Then, define $\mathcal{N}={\mathbb{N}_{>0}}$ and, for all $(n,\delta)\in{\mathbb{N}_{>0}}\times(0,1)$,
@@
    Assumption D.3 holds, $R$ is of trace class, and $\eta$ is $R$-subgaussian. Then, define $\mathcal{N}=\{N_{0}\}$ for some $N_{0}\in{\mathbb{N}_{>0}}$ and, for all $(n,\delta)\in{\mathbb{N}_{>0}}\times(0,1)$,
@@
Let $\delta\in(0,1)$. In case 2, it follows from Lemma D.4 that it holds with probability at least $1-\delta$ that
@@
where we applied the lemma with $A=(\mathrm{id}_{\mathcal{H}}+V_{N_{0}})^{-1/2}M_{N_{0}}^{\star}$, yielding $T=T_{N_{0},\lambda}=A\cdot(R\otimes\mathrm{id}_{\mathbb{R}^{N_{0}}})A^{\star}$ and. In case 1, it follows from Corollary D.20 that it holds with probability at least $1-\delta$ that, for all $N\in{\mathbb{N}_{>0}}$,
@@
The result then immediately follows from Lemma D.2, (20), and (22), noting that $\beta_{\lambda}(N,\delta)=S+\frac{1}{\sqrt{\lambda}}\gamma_{N,\lambda}(\delta)$. ∎
D.5 Proof of Theorem 4.3
@@
The result follows immediately by applying Theorem D.23 with $\mathcal{H}=\mathcal{H}_{K}$, $L_{n}=K(\cdot,X_{n})^{\star}$, and $\eta_{n}=y_{n}-\mathbb{E}[p](X_{n})$, for all $n\in{\mathbb{N}_{>0}}$. Indeed, for these choices, the optimization problem (15) is equivalent to that in (4), since for all $h\in\mathcal{H}_{K}$, $g\in\mathcal{G}$, and $n\in{\mathbb{N}_{>0}}$
@@
showing that the evaluation operator in $X_{n}$ is $K(\cdot,X_{n})^{\star}$. Furthermore, the process $\eta$ satisfies the appropriate assumption D.3 or D.5 depending on the considered case, since $D$ is assumed to be a process of transition pairs of $p$, or a process of independent transition pairs of $p$. In case 2, the assumptions of case 2 in Theorem D.23 are satisfied, and the result follows. Similarly, In case 1, the assumptions of case 1 in Theorem D.23 are satisfied by Proposition D.12, and the result follows as well. ∎
Appendix E Proof of Theorem 4.2
@@
$=\mathbb{P}\left[\exists(n_{1},n_{2},x)\in\mathcal{N}_{1}\times\mathcal{N}_{2}\times\mathcal{S}:\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x)\land\left\lVert f_{D^{(1)}_{:n_{1}}}^{(1)}(x)-f_{D^{(2)}_{:n_{2}}}^{(2)}(x)\right\rVert>\sum_{i=1}^{2}B_{i}\left(D^{(i)}_{:n},x\right)\right]$
@@
Appendix F Experiment details
@@
F.1 Illustrative example
Figure 1 illustrates our test in a simple example with one-dimensional inputs and outputs. We use a Gaussian kernel with bandwidth $\gamma^{2}=0.25$ on $\mathcal{X}=[-1,1]$ and the inhomogeneous linear kernel on $\mathcal{Z}=\mathbb{R}$. We sample two unit-norm mean functions $f_{1},f_{2}\in\mathcal{H}_{k}$, from which we collect data sets $D_{i}=\{(x^{(i)}_{j},f_{i}(x^{(i)}_{j})+\epsilon^{(i)}_{j})\}_{j=1}^{N}$, $i\in\{1,2\}$, of size $N=25$. The covariates $x^{(i)}_{j}$ are sampled uniformly from $\mathcal{X}$ and $\epsilon^{(i)}_{j}\sim\mathcal{N}(0,0.05^{2})$, such that the Markov kernel corresponding to each data set is $p_{i}(\cdot,x)=\mathcal{N}(\cdot\mid f_{i}(x),0.05^{2})$, where $\mathcal{N}(\cdot\mid\mu,s^{2})$ is the Gaussian measure on $\mathbb{R}$ with mean $\mu$ and variance $s^{2}$. Consequently, $H_{0}(x,p_{1},p_{2})$ is equivalent to $f_{1}(x)=f_{2}(x)$. We apply our test with analytical thresholds (Figure 1, left) and bootstrapped thresholds (Figure 1, right). For the analytical thresholds, we use the ground truth RKHS function norm (that is, $1$) and the Gaussian noise standard deviation for the corresponding upper bounds on these quantities in Theorem 4.3. Table 1 reports any additional hyperparameter values.
Table 1: Hyperparameters used in generating Figure 1.
Parameter Value
Regularization $\lambda=0.01$
Bootstrap resamples $1000$
F.2 Empirical error rates
We evaluate our test in well-specified settings by generating two data sets, $D_{1}$ and $D_{2}$, each containing $N\in\mathbb{N}$ transition pairs of the form
$D_{i}=\{(x^{(i)}_{j},f_{i}(x^{(i)}_{j})+\epsilon_{j}^{(i)})\}_{j=1}^{N},\quad i\in\{1,2\}.$ (32)
Here, the inputs $x^{(i)}_{j}$ are sampled according to some distribution over $\mathcal{X}$ and the terms $\epsilon^{(i)}_{j}$ are i.i.d. noise. We choose each $f_{i}$ and the noise distribution so that the corresponding Markov kernel
$p_{i}(\cdot,x^{(i)}_{j})=\mathbb{P}[f_{i}(x^{(i)}_{j})+\epsilon^{(i)}_{j}\in\cdot]$ (33)
admits a mean embedding $\mathbb{E}[p_{i}]$ in the vector-valued RKHS of a known scalar input kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and a known scalar output kernel $\kappa:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$. In particular, we choose $f_{i}$ directly from the RKHS $\mathcal{H}_{k}$ of $k$. In all our experiments in this section, we use a Gaussian kernel
@@ F.2 Empirical error rates
Choice of mean functions
We sample mean functions $f_{i}\in\mathcal{H}_{k}$ randomly and aggregate experiment results over many independent random choices of $f_{i}$. Specifically, as in [41], each $f_{i}$ is sampled by first sampling $m$ points $x_{1},...,x_{m}$ uniformly at random from $\mathcal{X}$ and then sampling $f_{i}$ uniformly from
$\left\{f=\sum_{j=1}^{m}\alpha_{j}k(\cdot,x_{j})\mid\alpha_{1},...,\alpha_{j}\in\mathbb{R},\|f\|_{\mathcal{H}_{k}}=r\right\}\subseteq\mathcal{H}_{k},$ (35)
@@
We quantify both the type $\mathrm{I}$ and type $\mathrm{II}$ error rates through the test’s empirical positive rate — the proportion of data set pairs on which at least one covariate triggers rejection in the region of interest — under two data-generation regimes. Under the global null ($H_{0}(x)$ holds for all $x\in\mathcal{X}$), this positive rate directly estimates the type $\mathrm{I}$ error. Under an alternative ($H_{0}(x)$ fails for some $x\in\mathcal{X}$), the same quantity measures the test’s power, from which we comptue the type $\mathrm{II}$ error as one minus that power.
@@
as the empirical positive rate. Repeating this under both null and alternative regimes while varying the significance level $\alpha$ and other parameters, we obtain detailed error-rate curves that reveal how closely the observed type $\mathrm{I}$ error matches its nominal level and how the test power depends on the various parameters.
Enforcing either regime requires appropriate choices of the mean functions $f_{1}$ and $f_{2}$, and noise distributions. For instance, to enforce the global null, it is sufficient (but depending on the kernel, not strictly necessary) to generate both data sets from identical mean functions $f_{1}=f_{2}$ with identical noise distribution. For the alternative, we can introduce a controlled discrepancy, such as choosing $f_{2}=f_{1}+h$ for some function $h\in\mathcal{H}_{k}$, or altering the noise distribution in the case where we introduce a kernel $\kappa$ to test for other properties such as higher moments (cf. Appendix A). Because the exact construction depends on the kernel $\kappa$ and the specific experiment, we defer those design details to the individual experiment descriptions that follow.
Finally, we always compute positive rates from $T=100$ independently sampled data sets and report and report curves that are averaged over $100$ independent experiment runs, together with 2.5% and 97.5% quantiles taken over the experiment runs.
F.2.1 Global and local sensitivity
Figure 2 presents the empirical positive rate of our test in three different settings: (1) when the $H_{0}$ is enforced enforced everwhere on $\mathcal{X}$ (Figure 2, left); (2) when $H_{0}$ is violated at frequently sampled covariates (Figure 2, middle); and (3) when $H_{0}$ is violated at rarely sampled covariates (Figure 2, right). In each scenario, we generate $N=100$ covariates with noisy responses $D_{i}=\{(x^{(i)}_{j},f_{i}(x^{(i)}_{j})+\epsilon^{(i)}_{j})\}_{j=1}^{N}$, $i\in\{1,2\}$, with i.i.d. Gaussian noise $\epsilon_{j}^{(i)}\sim\mathcal{N}(0,\sigma^{2})$. We use the inhomogeneous linear kernel for the kernel $\kappa$ on $\mathcal{Z}=\mathbb{R}$. Consequently, $H_{0}(x)$ is equivalent to $f_{1}(x)=f_{2}(x)$. In the following, we describe the details of each scenario.
@@ F.2.1 Global and local sensitivity
We draw a unit-norm function $f_{1}\in\mathcal{H}_{k}$ at random, set $f_{2}=f_{1}$, and sample covariates uniformly from $\mathcal{X}$. Consequently, $p_{1}=p_{2}$ such that every rejection is a false positive, and the empirical positive rate estimates our type $\mathrm{I}$ error.
@@
We again sample a unit-norm function $f_{1}$ randomly and independently draw a perturbation $h\in\mathcal{H}_{k}$. Setting $f_{2}=f_{1}+h$ creates a global discrepancy between $p_{1}$ and $p_{2}$ whose magnitude is controlled by the parameter $\xi$. Covariates remain sampled uniformly. In this setting, a rejection is taken to be a true positive, neglecting the edge case where $h(x)=0$ incidentally causes $f_{1}(x)=f_{2}(x)$. We then take the empirical positive rate as the true positive rate, translating directly to the type $\mathrm{II}$ error.
@@
We sample $f_{1}\in\mathcal{H}_{k}$ randomly and set $f_{2}=f_{1}+k(\cdot,0)$. Then, $p_{1}(\cdot,x)$ and $p_{2}(\cdot,x)$ are approximately equal for $x\in\mathcal{X}_{\mathrm{same}}\coloneqq\{x\in\mathcal{X}\mid f_{2}(x)<10^{-2}\}$, but differ significantly for $x\in\mathcal{X}_{\mathrm{diff}}=\mathcal{X}\setminus\mathcal{X}_{\mathrm{same}}$. To control how often this anomalous region $\mathcal{X}_{\mathrm{diff}}$ is sampled, we now sample covariates from a mixture; namely from $\mathcal{X}_{\mathrm{diff}}$ with probability $\theta$, and from $\mathcal{X}_{\mathrm{same}}$ with probability $1-\theta$. Varying $\theta$ then controls the difficulty of detecting a violation, and the empirical positive rate again reflects test power under this more challenging alternative.
@@
We compare our test to the conditional two-sample procedure of Hu and Lei, which casts hypothesis testing as a weighted conformal-prediction problem and builds a rank-sum statistic from estimated density ratios. In their framework, one classifier is trained on covariates alone to learn their marginal density ratios, and a second on the joint covariate–response pairs to learn the conditional density ratios. Because the baseline’s performance hinges on the accuracy of the density ratio learning method, we give it oracle-style information in each scenario: Firstly, we inform the method in all scenarios that the marginal densities are identical between the two samples. Secondly, in Scenarios 1 and 2, we supply the true Gaussian noise distribution and restrict density ratio estimation from the Gaussian parametric form to estimating only conditional means via kernel ridge regression (as in our test). In Scenario 3, we go one step further and provide the baseline with the ground truth conditional density ratios. This alignment makes outcomes directly comparable and focuses the comparison on the heart of each test, namely, global testing using one aggregated statistic versus our local, covariate-specific decisions.
We report any additional hyperparameter values in Table 2 for Scenarios 1 and 2, and in Table 3 for Scenario 3, respectively. The experiments were conducted on an Intel Xeon 8468 Sapphire CPU, using 10 GB of RAM and 50 core hours.

### v2

@@ E.4 Main result
Theorem E.23.
Under the notations of Section E.1, let $R\in\mathcal{L}_{\mathrm{b}}(\mathcal{G})$ be self-adjoint, positive semi-definite, and $S\in\mathbb{R}_{>0}$. Assume that $\lVert h^{\star}\rVert_{\mathcal{H}}\leq S$, and assume one of the following and define $\beta_{\lambda}$ and $\mathcal{N}$ accordingly:
@@
    Assumptions E.5 and E.13 hold, $R=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$ for some $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, and $\eta$ is $R$-subgaussian conditionally on $\mathcal{F}$. Then, define $\mathcal{N}={\mathbb{N}_{>0}}$ and, for all $(n,\delta)\in{\mathbb{N}_{>0}}\times(0,1)$,
@@
    Assumption E.3 holds, $R$ is of trace class, and $\eta$ is $R$-subgaussian. Then, define $\mathcal{N}=\{N_{0}\}$ for some $N_{0}\in{\mathbb{N}_{>0}}$ and, for all $(n,\delta)\in{\mathbb{N}_{>0}}\times(0,1)$,
@@
Let $\delta\in(0,1)$. In case 2, it follows from Lemma E.4 that it holds with probability at least $1-\delta$ that
@@
where we applied the lemma with $A=(\mathrm{id}_{\mathcal{H}}+V_{N_{0}})^{-1/2}M_{N_{0}}^{\star}$, yielding $T=T_{N_{0},\lambda}=A\cdot(R\otimes\mathrm{id}_{\mathbb{R}^{N_{0}}})A^{\star}$ and. In case 1, it follows from Corollary E.20 that it holds with probability at least $1-\delta$ that, for all $N\in{\mathbb{N}_{>0}}$,
@@
The result then immediately follows from Lemma E.2, (20), and (22), noting that $\beta_{\lambda}(N,\delta)=S+\frac{1}{\sqrt{\lambda}}\gamma_{N,\lambda}(\delta)$. ∎
E.5 Proof of Theorem 4.3
@@
The result follows immediately by applying Theorem E.23 with $\mathcal{H}=\mathcal{H}_{K}$, $L_{n}=K(\cdot,X_{n})^{\star}$, and $\eta_{n}=y_{n}-\mathbb{E}[p](X_{n})$, for all $n\in{\mathbb{N}_{>0}}$. Indeed, for these choices, the optimization problem (15) is equivalent to that in (4), since for all $h\in\mathcal{H}_{K}$, $g\in\mathcal{G}$, and $n\in{\mathbb{N}_{>0}}$
@@
showing that the evaluation operator in $X_{n}$ is $K(\cdot,X_{n})^{\star}$. Furthermore, the process $\eta$ satisfies the appropriate assumption E.3 or E.5 depending on the considered case, since $D$ is assumed to be a process of transition pairs of $p$, or a process of independent transition pairs of $p$. In case 2, the assumptions of case 2 in Theorem E.23 are satisfied, and the result follows. Similarly, In case 1, the assumptions of case 1 in Theorem E.23 are satisfied by Proposition E.12, and the result follows as well. ∎
Appendix F Proof of Theorem 4.2
@@
$=\mathbb{P}\left[\exists(n_{1},n_{2},x)\in\mathcal{N}_{1}\times\mathcal{N}_{2}\times\mathcal{S}:\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x)\right$.
$\qquad\qquad\left.\land\left\lVert f_{D^{(1)}_{:n_{1}}}^{(1)}(x)-f_{D^{(2)}_{:n_{2}}}^{(2)}(x)\right\rVert>\sum_{i=1}^{2}B_{i}\left(D^{(i)}_{:n},x\right)\right]$
@@
Appendix G Experiment details
@@
Unless stated, all our experiments with bootstrapped test thresholds use the naive resampling scheme outlined in Appendix B.
G.1 Illustrative example (Figure 1)
Figure 1 illustrates our test on a simple example with one-dimensional inputs and outputs. We use a Gaussian kernel on $\mathcal{X}=[-1,1]$ and the inhomogeneous linear kernel on $\mathcal{Z}=\mathbb{R}$. We pick two mean functions $f_{1},f_{2}\in\mathcal{H}_{k}$, from which we collect data sets $D_{i}=\{(x^{(i)}_{j},f_{i}(x^{(i)}_{j})+\epsilon^{(i)}_{j})\}_{j=1}^{n}$, $i\in\{1,2\}$. The covariates $x^{(i)}_{j}$ are sampled uniformly from $\mathcal{X}$ and $\epsilon^{(i)}_{j}\sim\mathcal{N}(\cdot\mid 0,s^{2})$, such that the Markov kernel corresponding to each data set is $p_{i}(\cdot,x)=\mathcal{N}(\cdot\mid f_{i}(x),s^{2})$, where $\mathcal{N}(\cdot\mid\mu,s^{2})$ is the Gaussian measure on $\mathbb{R}$ with mean $\mu$ and variance $s^{2}$. Consequently, $H_{0}(x,p_{1},p_{2})$ is equivalent to $f_{1}(x)=f_{2}(x)$.
We apply our test with analytical thresholds (Figure 1, left) and bootstrapped thresholds (Figure 1, right). For the analytical thresholds, we use the ground truth RKHS function norm (that is, $1$) and the Gaussian noise standard deviation for the corresponding upper bounds on these quantities in Theorem 4.3.
Table 2 reports the hyperparameters for this experiment.
G.2 Empirical error rates
This section provides implementation and design details for experiments evaluating the empirical type $\mathrm{I}$ and type $\mathrm{II}$ errors of our tests in controlled, well-specified settings. All experiments in this section were conducted on an Intel Xeon 8468 Sapphire CPU, using 10 GB of RAM.
G.2.1 General setup
Data generation
We evaluate our test by generating two data sets, $D_{1}$ and $D_{2}$, each containing $n\in\mathbb{N}$ transition pairs of the form
$D_{i}=\{(x^{(i)}_{j},f_{i}(x^{(i)}_{j})+\epsilon_{j}^{(i)})\}_{j=1}^{n},\quad i\in\{1,2\}.$
Here, the inputs $x^{(i)}_{j}$ are sampled (uniformly, unless stated otherwise) from $\mathcal{X}$, $\epsilon^{(i)}_{j}$ are i.i.d. noise, and $f_{i}$ is a random unit-norm element of $\mathcal{H}_{k}$ whose sampling is detailed below. In particular, we choose $f_{i}$ directly from the RKHS $\mathcal{H}_{k}$ of $k$. In all our experiments in this section, we use a Gaussian kernel
@@ G.2.1 General setup
Remark G.1.
In the case where $\kappa$ is the linear kernel, the procedure above ensures that $\mathbb{E}[p_{i}]\in\mathcal{H}_{K}$ with $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$. This guarantee is lost when $\kappa$ is a more complex kernel. While this means that the simulations may use the test in contexts not covered by the theory, this is precisely the interest of the bootstrapping schemes we propose: the test can still be applied, and its performance evaluated empirically.
Sampling of mean functions
Following Fiedler et al. [25], we sample mean functions $f_{i}\in\mathcal{H}_{k}$ randomly by first sampling $m$ points $x_{1},...,x_{m}$ uniformly at random from $\mathcal{X}$ and then sampling $f_{i}$ uniformly from the sphere
$\left\{f=\sum_{j=1}^{m}\alpha_{j}k(\cdot,x_{j})\bigm|\alpha_{1},...,\alpha_{m}\in\mathbb{R},\|f\|_{\mathcal{H}_{k}}=r\right\}\subseteq\mathcal{H}_{k},$ (33)
@@
We quantify both the type $\mathrm{I}$ and type $\mathrm{II}$ error rates through the test’s empirical positive rate — the proportion of data set pairs on which at least one covariate triggers rejection in the region of interest — under two data-generation regimes. Under the global null ($H_{0}(x,p_{1},p_{2})$ holds for all $x\in\mathcal{X}$), this positive rate directly estimates the type $\mathrm{I}$ error. Under an alternative ($H_{0}(x,p_{1},p_{2})$ fails for some $x\in\mathcal{X}$), the same quantity measures the test’s power, from which we compute the type $\mathrm{II}$ error as one minus that power.
@@
as the empirical positive rate.
Repeating this under the null and alternative regimes while varying the significance level $\alpha$ and other parameters, we obtain detailed error-rate curves that reveal how closely the observed type $\mathrm{I}$ error matches its nominal level and how the test power depends on the various parameters. Enforcing either regime requires appropriate choices of the mean functions $f_{1}$ and $f_{2}$, and noise distributions. For instance, to enforce the global null, it is sufficient (but depending on the kernel $\kappa$, not strictly necessary) to generate both data sets from identical mean functions $f_{1}=f_{2}$ with identical noise distribution. For the alternative, we can introduce a controlled discrepancy, such as choosing $f_{2}=f_{1}+h$ for some function $h\in\mathcal{H}_{k}$, or alter the noise distribution in the case where we introduce a kernel $\kappa$ to test for other properties such as higher moments. Because the exact construction for enforcing null and alternative hypotheses depends on the specific experiment setting, we defer those specific design details to the individual experiment descriptions that follow.
Finally, we always compute positive rates from $T=100$ independently sampled data sets and report curves that are averaged over $100$ independent experiment runs (i.e., samples of mean function pairs), together with 2.5% and 97.5% quantiles taken over the experiment runs.
G.2.2 Global and local sensitivity (Figure 2)
Figure 2 presents the empirical positive rate of our test in three different scenarios: (1) when the $H_{0}$ is enforced enforced everywhere on $\mathcal{X}$ (Figure 2, left); (2) when $H_{0}$ is violated at frequently sampled covariates (Figure 2, middle); and (3) when $H_{0}$ is violated at rarely sampled covariates (Figure 2, right). In all three scenarios, we use the inhomogeneous linear kernel for $\kappa$ and Gaussian noise $\mathcal{N}(0,s^{2})$ such that $H_{0}(x)$ holds if and only if $f_{1}(x)=f_{2}(x)$. In the following, we describe how each scenario enforces its corresponding hypothesis.
@@ G.2.2 Global and local sensitivity (Figure 2)
We pick a unit-norm function $f_{1}\in\mathcal{H}_{k}$ and set $f_{2}=f_{1}$. Consequently, $p_{1}=p_{2}$ such that every rejection is a false positive, and the empirical positive rate estimates our type $\mathrm{I}$ error.
@@
We pick a unit-norm function $f_{1}\in\mathcal{H}_{k}$, and independently pick a perturbation $h\in\mathcal{H}_{k}$ with norm $\xi$. By setting $f_{2}=f_{1}+h$, $f_{1}$ and $f_{2}$ differ almost everywhere almost surely, thereby creating a global discrepancy between $p_{1}$ and $p_{2}$ whose magnitude is controlled by $\xi$. In this setting, a rejection is taken to be a true positive, neglecting the edge case where $h(x)=0$ incidentally causes $f_{1}(x)=f_{2}(x)$. We then take the empirical positive rate as the true positive rate, translating directly to our type $\mathrm{II}$ error.
@@
We pick a unit-norm function $f_{1}\in\mathcal{H}_{k}$ and set $f_{2}=f_{1}+k(\cdot,0)$. Then, $p_{1}(\cdot,x)$ and $p_{2}(\cdot,x)$ are approximately equal for $x\in\mathcal{X}_{\mathrm{same}}\coloneqq\{x\in\mathcal{X}\mid k(\cdot,x)<10^{-2}\}$, but differ significantly for $x\in\mathcal{X}_{\mathrm{diff}}=\mathcal{X}\setminus\mathcal{X}_{\mathrm{same}}$. The value $10^{-2}$ was chosen as an order of magnitude at which it is unrealistic to expect detecting the difference from the considered data set sizes. To control how often this anomalous region $\mathcal{X}_{\mathrm{diff}}$ is sampled, we now sample covariates from a mixture; namely from $\mathcal{X}_{\mathrm{diff}}$ with probability $\theta$, and from $\mathcal{X}_{\mathrm{same}}$ with probability $1-\theta$. Varying $\theta$ then controls the difficulty of detecting a violation, and the empirical positive rate again translates to our type $\mathrm{II}$ error under this more challenging alternative.
@@
We compare our test to the conditional two‐sample procedure of Hu and Lei [5], which casts hypothesis testing as a weighted conformal‐prediction problem and builds a rank‐sum statistic from estimated density ratios. In their framework, one classifier is trained on covariates alone to learn their marginal density ratios, and a second on the joint covariate–response pairs to learn the conditional density ratios. Because the baseline’s performance hinges on the accuracy of the density ratio learning method, we provide it with oracle-style information in each scenario. Firstly, we inform the method in all scenarios that the marginal densities are identical between the two samples. Secondly, in Scenarios 1 and 2, we supply the true Gaussian noise distribution and restrict the conditional density-ratio model to the Gaussian parametric family with known variance, so the learner only estimates conditional means via KRR (as in our test). In Scenario 3, we go one step further and provide the baseline with the ground truth conditional density ratios. This alignment makes outcomes directly comparable and focuses the comparison on the heart of each test, namely, global testing using one aggregated statistic versus our local, covariate-specific decisions.
Table 3 reports the hyperparameters for Scenarios 1 and 2, and Table 4 those for Scenario 3. The experiments used 50 core hours.
G.2.3 Comparison of higher-order moments (Figure 4)
Figure 4 shows the positive rate of our test for different kernels $\kappa$ on $\mathcal{Z}=\mathbb{R}$ when the conditional means coincide, but the conditional distributions still differ. We pick a unit-norm mean function $f$ and set $f_{1}=f_{2}=f$, such that both data sets $D_{1}$ and $D_{2}$ in each pair are generated from the same mean function. However, $D_{1}$ and $D_{2}$ differ in the noise applied to the observations. The observations in $D_{1}$ are corrupted with Gaussian noise $\mathcal{N}(0,s^{2})$. In contrast, the observations in $D_{2}$ are corrupted with noise sampled from an equally weighted mixture of $\mathcal{N}(-\mu,s^{2})$ and $\mathcal{N}(\mu,s^{2})$. For this data-generating process, the conditional means coincide while the conditional distributions differ. We compute the empirical positive rates in this setting for (i) a linear kernel; and (ii) a Gaussian kernel on $\mathcal{Z}$.
Table 6 reports the hyperparameters for this experiment. The experiment used 20 core hours.
G.2.4 Influence of using an overly rich kernel (Figure 5)
Figure 5 shows the type $\mathrm{I}$ and $\mathrm{II}$ errors of our test for different Gaussian and polynomial kernels $\kappa$ on $\mathcal{Z}=\mathbb{R}$. For the type $\mathrm{I}$ error, we pick a single unit-norm mean function $f=f_{1}=f_{2}$ and report the empirical positive rate. For the type $\mathrm{II}$ error, we pick two mean functions $f_{1}$ and $f_{2}$ independently. In either case, we apply the same Gaussian noise $\mathcal{N}(0,s^{2})$ to both data sets.
Table 7 reports the hyperparameters for this experiment. The experiment used 15 core hours.
G.2.5 Bootstrapping schemes (Figures 6 and 7)
Figure 6 shows the type $\mathrm{I}$ and $\mathrm{II}$ errors of our test for different bootstrapping schemes when varying the data set size and regularization of the KRR. For the type $\mathrm{I}$ error, we pick a single unit-norm mean function $f=f_{1}=f_{2}$ and report the empirical positive rate. For the type $\mathrm{II}$ error, we pick two mean functions $f_{1}$ and $f_{2}$ independently. We use a linear kernel $\kappa$ on $\mathcal{Z}$ and apply the same Gaussian noise $\mathcal{N}(0,s^{2})$ to both data sets. In Figure 6(a), we fix the regularization $\lambda=0.5$ and vary the data set size $n$. In Figure 6(b), we fix $n=50$ and vary $\lambda$.
Figure 7 reports type $\mathrm{I}$ and $\mathrm{II}$ errors for analytical and bootstrapped test thresholds, using the same setup as before but fixing both $n=100$ and $\lambda=0.25$.
Tables 8 and 9 report the hyperparameters for these experiments. The experiments used 30 core hours.
G.3 Process monitoring (Figure 3)
Figure 3 shows the ratio between the test statistic and threshold averaged over the sliding window following the trajectory of a perturbed linear dynamical system (left), and the average value of $\sigma_{D,\lambda}(x)$ over the window when $D$ is the reference data set. We detail this experiment here.
We set $\mathcal{X}=\mathbb{R}^{d}$, $d\in{\mathbb{N}_{>0}}$, and choose the linear operator-valued kernel $K(x,x^{\prime})=\langle x,x^{\prime}\rangle_{\mathbb{R}^{d}}\cdot\mathrm{id}_{\mathbb{R}^{d}}$, as we wish to compare linear functions mapping $\mathbb{R}^{d}$ to itself. We randomly sample a matrix $A\in\mathrm{O}(d)$, the group of orthonormal matrices in dimension $d$, pick a fixed initial state $X_{0}=\frac{1}{\sqrt{d}}[1\cdots 1]^{\top}\in\mathbb{R}^{d}$, and generate a single trajectory $(X_{n})_{n\in\left[N\right]}$ as $X_{n+1}=AX_{n}+\epsilon_{n}$, where $\epsilon_{n}\sim\mathcal{N}(0,s^{2}I_{d})$ is i.i.d. Gaussian noise. We enforce that $A$ is orthonormal so the trajectory neither shrinks nor explodes exponentially. Then, the observations $Y=(Y_{n})_{n\in{\mathbb{N}_{>0}}}$ are taken as $Y_{n}=X_{n+1}$. It is immediate to verify that $((X_{n},Y_{n}))_{n\in{\mathbb{N}_{>0}}}$ is a process of transition pairs of the Markov kernel $p(\cdot,x)=\mathcal{N}(\cdot\mid Ax,s^{2}I_{d})$, where $\mathcal{N}(\cdot\mid\mu,s^{2}I_{d})$ is the Gaussian measure on $\mathbb{R}^{d}$ with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $s^{2}I_{d}\in\mathbb{R}^{d\times d}$.
Using this procedure for $5$ independent system trajectories of length $400$, we collect a data set of $n_{\mathrm{ref}}=2000$ data points, which we call the reference data set $D$. We then follow the trajectory with a sliding window of size $n_{\mathrm{win}}=50$ data points, and perform the test at every step on the points in the sliding window. After $200$ steps, we introduce a change in the matrix $A$, replacing it with the matrix $A^{\prime}=A\exp(\frac{\xi}{\|H\|_{1}}H)$. Here $H=\frac{1}{2}(B-B^{\top})$ is antisymmetric and the entries of $B$ are sampled i.i.d. from $\mathcal{N}(0,1)$, such that $\mathrm{dist}_{\mathrm{O}(d)}(A,A^{\prime})=\xi$, where $\mathrm{dist}_{\mathrm{O}(d)}(A,A^{\prime})=\|\log(A^{\top}A^{\prime})\|_{\mathrm{HS}}$ is the usual geodesic distance on $\mathrm{O}(d)$. This construction ensures that $A^{\prime}$ is orthonormal as well, and controls the distance between $A$ and $A^{\prime}$ in a way consistent with the geometry of $\mathrm{O}(d)$.
Figure 3 (left) shows the result of this procedure, for varied values of $d$ and of $\xi$ ($\xi$ is only varied with $d=16$), and averaged over $100$ independent choices of the data sets and $50$ independent choices of the matrix $A$. We observe that the change is detected rapidly, even before the sliding window is entirely filled with data from the new dynamics. Notably, higher dimensions require higher disturbances to be successfully detected with a fixed amount of data. Figure 3 (right) suggests a reason for this: the variance of the reference data set at the current state increases after change, indicating that the dynamics with $A^{\prime}$ bring the trajectory into regions unexplored when following the dynamics with $A$. In other words, the lack of power in high dimensions seems to result from a lack of local data, as it is “easier” to go in unsampled regions by changing the dynamics.
Table 5 presents the hyperparameters for this experiment. The experiment used 2000 core hours.
Table 2: Hyperparameters used in generating Figure 1.
Parameter Value
Input set $\mathcal{X}=[-1,1]$
Input kernel bandwidth $\gamma^{2}=0.25$
Data set size $n=25$
Noise variance $s^{2}=0.05^{2}$
Regularization $\lambda=0.01$
Bootstrap resamples $M=1000$